3x + 2y =16
2x + y =9

Answer by checkley71(8403)

3x+2y=16 & 2x-y=9 mult eq#2 by 2 and subtract thus
4x+2y=18
----------
-x=-2 or x=2 or 3*2+2y=16 or 6+2y=16 or 2y=10 y=5